# Turbocharging constraints on dark matter substructure through a synthesis of strong lensing flux ratios and extended lensed arcs

### v2

@@ 3.1 Multi-plane lens modeling
    We evaluate the surface brightness of the source light in the souce plane, and then cast this light back to the image plane to produce a lensed image. We then evaluate the likelihood of the imaging data.
@@ 3.2 The decoupled multi-plane approximation for multi-plane lensing
Equation 9 increases the speed of lens mass and source light reconstruction with line-of-sight halos included in the lens model because it decouples deflections produced by halos behind the main deflector from deflections produced in the main lens plane. Thus, using Equation 9 requires only one full ray tracing calculation to the source plane per coordinate in the image plane, circumventing the repeated evaluation of step $iii$ for each new proposal of $\boldsymbol{\alpha_{\rm{macro}}}$. A lens mass and source light reconstruction with substructure performed according to this procedure takes less than ten minutes, corresponding to an increase in speed by a factor between 100-1000, depending on the number of line-of-sight halos behind the main deflector. We have added this decoupled multi-plane approximation for option use in lenstronomy  [^5] This notebook illustrates the functionality in combination with pyHalo..
@@ 3.3 Validity of the decoupled multi-plane approximation
We can check the validity of the approximation discussed in the previous section for predicting image positions and flux ratio statistics through comparisons with the statistics obtained by exact multi-plane ray tracing. First, using exact ray tracing techniques discussed by Gilman et al. (2019) we compute $p_{{\rm{eqn3}}}\left(\boldsymbol{O_{\rm{pos}}},\boldsymbol{O_{\rm{f}}}|\boldsymbol{q_{\rm{cdm}}}\right)$, the probability of observing a given set of flux ratios with subhalos and line-of-sight halos included in the lens model with properties as prescribed by CDM (represented by $\boldsymbol{q_{\rm{cdm}}}$). We then repeat this procedure using the ray tracing approximation discussed in the previous section to compute the likelihood $p_{{\rm{eqn9}}}\left(\boldsymbol{O_{\rm{pos}}},\boldsymbol{O_{\rm{f}}}|\boldsymbol{q_{\rm{cdm}}}\right)$. For details regarding the substructure models we use for these tests, we refer to Section 4.1. We perform these calculations using the image positions of a reference mock lens model created without substructure (hereafter, the “smooth model”). This reference model has lens (source) redshifts $z_{\rm{d}}=0.5$ ($z_{\rm{source}}=1.5$), and a mass profile parameterized as an elliptical power law with Einstein radius of $1\ \rm{arcsec}$, an axis ratio of $0.75$, a logarithmic mass profile slope of $-2.0$, and external shear strength of $0.09$.
Figure 3 shows the joint likelihoods for the three flux ratios of the mock lens, with the flux ratios of a reference lens model with the same lens model smooth lens model marked with the green points. The likelihood computed through exact ray tracing is shown in black, while the likelihood computed with the approximation introduced in the previous sub-section is represented in blue. The likelihoods peak at the smooth model flux ratio values, which is simply a statement that substructure introduces perturbations around the flux ratios predicted by the smooth lens model. As the likelihood contours appear indistinguishable, we can conclude that the approximation presented in the previous section does not introduce a detectable bias in the model-predicted image flux ratios for a CDM dark matter model.
Unlike the flux ratios and image positions, it is computationally intractable to compare our approximation with results obtained through exact ray tracing methods for reasons discussed in Section 3.1. Instead, we perform tests with simulated datasets to verify that we obtain unbiased inferences of dark matter substructure properties when using this approximation to reconstruct imaging data with substructure included in the lens model. We describe the details of these simulations in the following section.
@@ 4 Tests on simulated data
This section details the how we create simulated datasets with which to test the lens modeling methodology discussed in the previous section. Section 4.1 discusses the warm dark matter model on which we test the methodology. Section 4.2 discusses the properties of the simulated main deflectors we use in our simulations. Section 4.3 details the parameterization of the lens and source light models used to create and model the simulated datasets. Section 4.4 details the modeling assumptions related to the observing conditions and point spread function. Throughout this section we use $\mathcal{U}$ to represent a uniform prior and $\mathcal{G}$ to represent a Gaussian prior.
@@ 4.1 Modeling of dark matter subhalos and line-of-sight halos
with $a=2.3$, $b=0.8$, and $c=-1.0$ (Lovell, 2020), and $dN_{\rm{CDM}}/dm$ represents the (sub)halo mas function in CDM. We draw line-of-sight halos the Sheth-Tormen (Sheth et al., 2001) halo mass function, and generate subhalos from a mass function of the form
@@
We use a pivot scale $m_{0}=10^{8}M_{\odot}$, and a logarithmic profile slope $\alpha=1.9$ (Springel et al., 2008). We apply the same suppression function in Equation 10 to both the line-of-sight and subhalo mass functions.
@@
We create a sample of 25 lenses with a CDM ground truth using $\Sigma_{\rm{sub}}=0.05\ \rm{kpc^{-2}}$ and $m_{\rm{hm}}=0$. The chosen value of $\Sigma_{\rm{sub}}$ roughly corresponds to the amount of substructure inferred by Gilman et al. (2020) and is consistent with N-body to within $\mathcal{O}\left(10\right)$ factor (Fiacconi et al., 2016).
@@
Figure 9: The inference on the macromodel parameters in Mock 4. The black distribution shows the inference obtained from using only image positions and flux ratios to constrain the lens model.The blue distribution results from using image positions and imaging data to constrain the lens model, and the magenta results from using image positions, flux ratios, and imaging data. These distributions are marginalized over the properties of substructure in each system. From left, the x-axis labels correspond to the normalization of the main deflector mass profile $\theta_{\rm{E}}$, the axis ratio $q$, the position angle of the main deflector ellipticity $\phi_{\rm{q}}$, the external shear strength $\gamma_{\rm{ext}}$, the position angle of the external shear $\phi_{\rm{\gamma\rm{ext}}}$, the logarithmic profile slope of the main deflector $\gamma$, and the multipole moments $a_{3}$ and $a_{4}$. Contours enclose $68\%$ and $95\%$ confidence intervals, and the true lens model parameters for the mock lens are marked with the green crosshairs.
@@
Figure 10: The inference on the macromodel parameters in Mock 6 obtained from only the image positions and flux ratios (black), from the image positions and imaging data (blue), and from the image positions, flux ratios, and imaging data (magenta).
@@
Figure 11: The inference on the macromodel parameters in Mock 11 obtained from only the image positions and flux ratios (black), from the image positions and imaging data (blue), and from the image positions, flux ratios, and imaging data (magenta).
@@
Figure 12: The inference on the macromodel parameters in Mock 23 obtained from only the image positions and flux ratios (black), from the image positions and imaging data (blue), and from the image positions, flux ratios, and imaging data (magenta).
@@
Figure 13: The likelihood functions for the normalization of the subhalo mass function, $\Sigma_{\rm{sub}}$, and the turnover scale of a WDM mass function, $m_{\rm{hm}}$, for Mocks 4, 6, 11, and 23, computed following the down-selection on flux ratio and imaging data summary statistics as described in Section 2.2. The color scale indicates relative likelihood between different points in parameter space.
@@
Figure 14: Reconstructed lensed image (left), substructure convergence (center), and normalized residual map from the reconstructed imaging data of Mock $4$. The top two rows depict realizations accepted based matching the image positions, flux ratios, and imaging data. The bottom rows show examples of systems that match the flux ratios, but which we reject due to a poor fit to the imaging data. The green (black) numbers and curves show the true (model-predicted) flux ratios and critical curves, respectively.
@@
Figure 15: The same as Figure 14, but with four example realizations generated for Mock $6$. Simultaneously reproducing the flux ratios and imaging data in this system requires dark matter substructure near image A. As shown by the rejected lens model configuration in the third row, incorporating imaging data allows us to rule out lens model configurations that match the flux ratios through large-scale deformation of the lens mass profile, isolating the effects of substructure on these data.
@@
Figure 16: The same as Figure 14, but with four example realizations generated for Mock $11$. This system has flux ratios consistent with those predicted by a smooth lens model, so accepted realizations are more likely to have fewer halos. As seen in the bottom row, the information encoded by the imaging data allows us to rule out lens model configurations that match the flux ratios through a combination of substructure and large-scale deformation of the lens mass profile.
@@
Figure 17: The same as Figure 14, but with four example realizations generated for Mock $23$. Proposed lens models accepted for this system tend to have substructure perturbing image B.
@@ 4.2.1 Creating mock lens mass profiles
where $r$ is the separation in arcseconds from the mass centroid, the angle $\phi_{m}$ determines the orientation, and $a_{\rm{m(phys)}}$ is the amplitude of the convergence associated with the multipole perturbation. Oh et al. (in prep) further explore the implications of these multipole perturbations in the context of flux ratio analyses in strong lenses.
We base our implementation of the multipole perturbations on the observed properties of 847 elliptical galaxies presented by Hao et al. (2006). The shape of the iso-density contours inferred from the light (and assuming light traces mass) can be related to the physical amplitude of the perturbation by $a_{\rm{m(phys)}}=a_{\rm{m}}\times\frac{\theta_{E}}{\sqrt{q}}$, where $\theta_{\rm{E}}$ is the Einstein radius, $q$ is the axis ratio of the ellipse and $a_{m}$ represents the observable deviation from ellipticity of the light. The third ($m=3$) and fourth-order ($m=4$) moments make the dominant contribution to the multipole expansion of galaxy shapes (Bender et al., 1989). Hao et al. (2006) find $a_{\rm{3}}$ distributed as $\mathcal{G}\left(0.0,0.005\right)$ with values of $\phi_{3}$ uncorrelated with the position angle of the underlying mass profile. Hao et al. (2006) also measure $a_{4}\in\mathcal{G}\left(0.0,0.01\right)$ with an orientation that tends to align with the orientation of the underlying mass profile. When the $a_{4}$ orientation aligns with the orientation of the underlying ellipse the profile appears boxy ($a_{4}<0$) or disky ($a_{4}>0$). The observed amplitudes $a_{3}$ and $a_{4}$ have no apparent correlation.
Based on the findings by Hao et al. (2006), when creating mock deflectors we include the dominant $m=3$ and $m=4$ terms and use priors for their observed amplitudes $a_{3}\in\mathcal{G}\left(0.0,0.005\right)$ and $a_{4}\in\mathcal{G}\left(0.0,0.01\right)$. Based on the measurements by Hao et al. (2006) and Oh et al. (in prep), we enforce alignment between the $a_{4}$ orientation and the underlying EPL profile (producing boxy or disky contours), and sample $\phi_{3}\in\mathcal{U}\left(-\pi/6,\ \pi/6\right)$.
@@ 4.3.1 Creating mock lens and source light profiles
We parameterize the main deflector light as a circular Sérsic profile (Sérsic, 1963). The source light model includes two components: the lensed quasar and its host galaxy. For the quasar, we assume flux ratios measured from the warm dust region now observable with JWST (Nierenberg et al., 2023) with a physical size $\sim 1-10\ \rm{pc}$. To create the mocks, we model the quasar emission as a circular Gaussian with a full-width at half maximum sampled from a uniform prior $\mathcal{U}\left(1-10\right)\rm{pc}$. We assume a flux ratio measurement precision of $3\%$, and an astrometric precision in the relative quasar image positions of 5 milli-arcseconds, based on the recent measurements (Nierenberg et al., 2023).

### v5

@@ 3.1 Multi-plane lens modeling
    We evaluate the surface brightness of the source light in the souce plane, and then cast this light back to the image plane to produce a lensed image. We then evaluate the imaging data likelihood.
@@ 3.2 The decoupled multi-plane approximation for multi-plane lensing
Equation 9 increases the speed of lens mass and source light reconstruction with line-of-sight halos included in the lens model because it decouples deflections produced by halos behind the main deflector from deflections produced in the main lens plane. Thus, using Equation 9 requires only one full ray tracing calculation to the source plane per coordinate in the image plane, circumventing the repeated evaluation of step $iii$ for each new proposal of $\boldsymbol{\alpha_{\rm{macro}}}$. A lens mass and source light reconstruction with substructure performed according to this procedure takes less than ten minutes, corresponding to an increase in speed by a factor between 100-1000, depending on the number of line-of-sight halos behind the main deflector. We have added this decoupled multi-plane approximation for optional use in lenstronomy  [^5] This notebook illustrates the functionality in combination with pyHalo..
@@ 3.3 Validity of the decoupled multi-plane approximation
We can check the validity of the approximation discussed in the previous section for predicting image positions and flux ratio statistics through comparisons with the statistics obtained by exact multi-plane ray tracing. First, using exact ray tracing techniques discussed by Gilman et al. (2019) we compute $p_{{\rm{eqn3}}}\left(\boldsymbol{O_{\rm{pos}}},\boldsymbol{O_{\rm{f}}}|\boldsymbol{q_{\rm{cdm}}}\right)$, the probability of observing a given set of flux ratios with subhalos and line-of-sight halos included in the lens model with properties as prescribed by CDM (represented by $\boldsymbol{q_{\rm{cdm}}}$). We then repeat this procedure using the ray tracing approximation discussed in the previous section to compute the likelihood $p_{{\rm{eqn9}}}\left(\boldsymbol{O_{\rm{pos}}},\boldsymbol{O_{\rm{f}}}|\boldsymbol{q_{\rm{cdm}}}\right)$. For details regarding the substructure models we use for these tests, we refer to Section 4.1. We perform these calculations using the image positions of a reference mock lens model created without substructure (hereafter, the “smooth model”). This reference model has a cross image configuration with lens (source) redshifts $z_{\rm{d}}=0.5$ ($z_{\rm{source}}=1.5$), and a mass profile parameterized as an elliptical power law with Einstein radius of $1\ \rm{arcsec}$, an axis ratio of $0.75$, a logarithmic mass profile slope of $-2.0$, and external shear strength of $0.09$. We have also performed a similar test for a fold image configuration.
Figure 3 shows the joint likelihoods for the three flux ratios of the mock lens, with the flux ratios of the reference smooth lens model marked with the green points. The likelihood computed through exact ray tracing is shown in black, while the likelihood computed with the approximation introduced in the previous sub-section is represented in blue. The likelihoods peak at the smooth model flux ratio values, which is simply a statement that substructure introduces perturbations around the flux ratios predicted by the smooth lens model. A dark matter inference with flux ratios requires the characterization of this multi-dimensional probability density. Dark matter substructure deforms the structure of this probability distribution in a distinct way from other sources of small-scale perturbation, such as multipole terms added to the main deflector mass profile.
[…]
Unlike the flux ratios and image positions, it is computationally intractable to compare our approximation with exact ray tracing when reconstructing imaging data for the reasons discussed in Section 3.1. Instead, we perform tests with simulated datasets to verify that we obtain unbiased inferences of dark matter substructure properties when using this approximation to reconstruct imaging data with substructure included in the lens model. We describe the details of these simulations in the following section.
@@ 4 Tests on simulated data
This section details the how we create simulated datasets with which to test the lens modeling methodology discussed in the previous section. Section 4.1 discusses the warm dark matter model on which we test the methodology. Section 4.2 discusses the properties of the simulated main deflectors we use in our simulations. Section 4.3 details the parameterization of the lens and source light models used to create and model the simulated datasets. Section 4.4 details the modeling assumptions related to the observing conditions and point spread function. Throughout this section we use $\mathcal{U}$ to represent a uniform prior and $\mathcal{N}$ to represent a Gaussian prior, not be confused with $\mathcal{N}$, the vector of nuisance parameters appearing in Equation 2.
@@ 4.1 Modeling of dark matter subhalos and line-of-sight halos
with $a=2.3$, $b=0.8$, and $c=-1.0$ (Lovell, 2020), and $dN_{\rm{CDM}}/dm$ represents the (sub)halo mas function in CDM. We draw line-of-sight halos from the Sheth-Tormen (Sheth et al., 2001) halo mass function, and generate subhalos from a mass function of the form
@@
We use a pivot scale $m_{0}=10^{8}M_{\odot}$, and a logarithmic profile slope $\alpha=1.9$ (Springel et al., 2008). The normalization is related to the projected mass in substructure in the range $10^{6}-10^{10}$ solar masses by $f_{\rm{proj}}=10^{7}\left(\Sigma_{\rm{sub}}/0.01\ \rm{kpc^{-2}}\right)\ \rm{M_{\odot}}\ \rm{kpc^{-2}}$, and N-body simulations of massive ellipticals predict typical values in the range $\sim 2-3\times 10^{7}M_{\odot}/\rm{kpc^{-2}}$ (e.g Fiacconi et al., 2016), although these numbers are uncertain due to various numerical uncertainties inherent to the simulations, such as the artificial disruption of subhalos and subhalo finders. To encompass a broad range of theoretical uncertainties associated with $\Sigma_{\rm{sub}}$, we use a log-uniform prior between $\log_{10}\Sigma_{\rm{sub}}=-2.5$ and $\log_{10}\Sigma_{\rm{sub}}=-1.0$. We apply the same suppression function in Equation 10 to both the line-of-sight and subhalo mass functions.
@@
We create two sets of simulated data with which to test our inference methodology. First, we create a sample of 25 lenses with a CDM ground truth using $\Sigma_{\rm{sub}}=0.05\ \rm{kpc^{-2}}$ and $m_{\rm{hm}}=0$. The chosen value of $\Sigma_{\rm{sub}}$ roughly corresponds to the amount of substructure inferred by Gilman et al. (2020) and is consistent with N-body to within an $\mathcal{O}\left(10\right)$ factor (Fiacconi et al., 2016). Second, we create a sample of 25 lenses with a WDM ground truth with $\Sigma_{\rm{sub}}=0.04\ \rm{kpc^{-2}}$ and $m_{\rm{hm}}=10^{7.5}M_{\odot}$. Both sets of mocks have the same main deflector mass models and background sources, but the image positions and flux ratios between them differ slightly due to the different populations of halos in the lens models.
@@
Figure 9: The inference on the macromodel parameters in Mock 4 in the CDM ground truth sample. The black distribution shows the inference obtained from using only image positions and flux ratios to constrain the lens model.The blue distribution results from using image positions and imaging data to constrain the lens model, and the magenta results from using image positions, flux ratios, and imaging data. These distributions are marginalized over the properties of substructure in each system. From left, the x-axis labels correspond to the normalization of the main deflector mass profile $\theta_{\rm{E}}$, the axis ratio $q$, the position angle of the main deflector ellipticity $\phi_{\rm{q}}$, the external shear strength $\gamma_{\rm{ext}}$, the position angle of the external shear $\phi_{\rm{\gamma\rm{ext}}}$, the logarithmic profile slope of the main deflector $\gamma$, and the multipole moments $a_{3}$ and $a_{4}$. Contours enclose $68\%$ and $95\%$ credible intervals, and the true lens model parameters for the mock lens are marked with the green crosshairs.
@@
Figure 10: The inference on the macromodel parameters in Mock 6 in the CDM ground truth sample obtained from only the image positions and flux ratios (black), from the image positions and imaging data (blue), and from the image positions, flux ratios, and imaging data (magenta).
@@
Figure 11: The inference on the macromodel parameters in Mock 11 in the CDM ground truth sample obtained from only the image positions and flux ratios (black), from the image positions and imaging data (blue), and from the image positions, flux ratios, and imaging data (magenta).
@@
Figure 12: The inference on the macromodel parameters in Mock 23 in the CDM ground truth sample obtained from only the image positions and flux ratios (black), from the image positions and imaging data (blue), and from the image positions, flux ratios, and imaging data (magenta).
@@
Figure 13: The likelihood functions for the normalization of the subhalo mass function, $\Sigma_{\rm{sub}}$, and the turnover scale of a WDM mass function, $m_{\rm{hm}}$, for Mocks 4, 6, 11, and 23 in the CDM ground truth sample. These likelihoods result from the down-selection on flux ratio and imaging data summary statistics described in Section 2.2. The color scale indicates relative likelihood between different points in parameter space.
@@
Figure 14: Reconstructed lensed image (left), substructure convergence (center), and normalized residual map from the reconstructed imaging data of Mock $4$ in the CDM ground truth sample. The top two rows depict realizations accepted based matching the image positions, flux ratios, and imaging data. The bottom rows show examples of systems that match the flux ratios, but which we reject due to a poor fit to the imaging data. The green (black) numbers and curves show the true (model-predicted) flux ratios and critical curves, respectively.
@@
Figure 15: The same as Figure 14, but with four example realizations generated for Mock $6$ in the CDM ground truth sample. Simultaneously reproducing the flux ratios and imaging data in this system requires dark matter substructure near image A. As shown by the rejected lens model configuration in the third row, incorporating imaging data allows us to rule out lens model configurations that match the flux ratios through large-scale deformation of the lens mass profile, isolating the effects of substructure on these data.
@@
Figure 16: The same as Figure 14, but with four example realizations generated for Mock $11$ in the CDM ground truth sample. This system has flux ratios consistent with those predicted by a smooth lens model, so accepted realizations are more likely to have fewer halos. As seen in the bottom row, the information encoded by the imaging data allows us to rule out lens model configurations that match the flux ratios through a combination of substructure and large-scale deformation of the lens mass profile.
@@
Figure 17: The same as Figure 14, but with four example realizations generated for Mock $23$ in the CDM ground truth sample. Proposed lens models accepted for this system tend to have substructure perturbing image B.
@@ 4.2.1 Creating mock lens mass profiles
where $r$ is the separation in arcseconds from the mass centroid, the angle $\phi_{m}$ determines the orientation, and $a_{\rm{m(phys)}}$ is the amplitude of the convergence associated with the multipole perturbation. Oh et al. (2024) further explore the implications of these multipole perturbations in the context of flux ratio analyses in strong lenses.
We base our implementation of the multipole perturbations on the observed properties of 847 elliptical galaxies presented by Hao et al. (2006). The shape of the iso-density contours inferred from the light (and assuming light traces mass) can be related to the physical amplitude of the perturbation by $a_{\rm{m(phys)}}=a_{\rm{m}}\times\frac{\theta_{E}}{\sqrt{q}}$, where $\theta_{\rm{E}}$ is the Einstein radius, $q$ is the axis ratio of the ellipse and $a_{m}$ represents the deviation from ellipticity of the light. The third ($m=3$) and fourth-order ($m=4$) moments make the dominant contribution to the multipole expansion of galaxy shapes (Bender et al., 1989). Hao et al. (2006) find $a_{\rm{3}}$ distributed as $\mathcal{N}\left(0.0,0.005\right)$ with values of $\phi_{3}$ uncorrelated with the position angle of the underlying mass profile. Hao et al. (2006) also measure $a_{4}\in\mathcal{N}\left(0.0,0.01\right)$ with an orientation that tends to align with the orientation of the underlying mass profile. When the $a_{4}$ orientation aligns with the orientation of the underlying ellipse the profile appears boxy ($a_{4}<0$) or disky ($a_{4}>0$). The observed amplitudes $a_{3}$ and $a_{4}$ have no apparent correlation.
Based on the findings by Hao et al. (2006), when creating mock deflectors we include the dominant $m=3$ and $m=4$ terms and use priors for their observed amplitudes $a_{3}\in\mathcal{N}\left(0.0,0.005\right)$ and $a_{4}\in\mathcal{N}\left(0.0,0.01\right)$. Based on the measurements by Hao et al. (2006) and (Oh et al., 2024), we enforce alignment between the $a_{4}$ orientation and the underlying EPL profile (producing boxy or disky contours), and sample $\phi_{3}\in\mathcal{U}\left(-\pi/6,\ \pi/6\right)$. For additional discussion regarding the role of multipole perturbations in dark matter inferences with quadruply-imaged quasars, we refer to Appendix B.
@@ 4.3.1 Creating mock lens and source light profiles
We parameterize the main deflector light as a circular Sérsic profile (Sérsic, 1963). The source light model includes two components: the lensed quasar and its host galaxy. For the quasar, we assume flux ratios measured from the warm dust region now observable with JWST (Nierenberg et al., 2024) with a physical size $\sim 1-10\ \rm{pc}$. To create the mocks, we model the quasar emission as a circular Gaussian with a full-width at half maximum sampled from a uniform prior $\mathcal{U}\left(1-10\right)\rm{pc}$. We assume a flux ratio measurement precision of $3\%$, and an astrometric precision in the relative quasar image positions of 5 milli-arcseconds, based on the recent measurements (Nierenberg et al., 2024).